# Stability of weighted minimal hypersurfaces under a lower $1$-weighted Ricci curvature bound

### v1

@@ 4.1. Non-existence
Let $(M,g,f)$ be a complete weighted manifold. We assume $\operatorname{Ric}_{f}^{1}>0$. Then there is no complete two-sided immersed stable $f$-minimal hypersurface $\Sigma$ in $M$ with $m_{f,\Sigma}(\Sigma)<+\infty$.
@@
Impera-Rimoldi [29] have presented non-existence results for stable $f$-minimal hypersurfaces satisfying volume growth conditions under the positivity of $\infty$-weighted Ricci curvature (see [29, Theorems A and B]). We show similar results for $1$-weighted Ricci curvature.
@@
The proof is done by contradiction. We possess the following relation of Brooks type between the bottom of the spectrum $\lambda_{1,\Sigma}$ of $\Delta_{f,\Sigma}$ and the volume entropy (cf. [7, Theorem 1], [29, Theorem A]):
@@ 2. $(2)$
We give a proof by contradiction. If $(1)$ holds, then [29, Proposition 6] together with Lemma 2.1 implies
@@
By the first condition in (4.2) and [29, Proposition 7], for every compact subset $\Omega$ in $\Sigma$, the bottom of the spectrum of the stability operator $\mathcal{L}_{f,\Sigma}$ over $\Sigma\setminus\Omega$ is negative. This contradicts with the finiteness of $f$-index and [29, Proposition 5]. $\Box$
@@ 4.2. Smooth compactness
Choi-Schoen [11] have established a smooth compactness of closed minimal surfaces in a compact three-dimensional manifold of positive Ricci curvature. Sharp [53] has extended their smooth compactness to the higher dimensional case. In the weighted case, such smooth compactness have been investigated by Li-Wei [37], Cheng-Mejia-Zhou [8], [9] and Barbosa-Sharp-Wei [4] under the positivity of $\infty$-weighted Ricci curvature (see also [12], [15] for self-shrinker). We prove smooth compactness under the positivity of $1$-weighted Ricci curvature.
We first show the following Frankel type estimate (cf. [20, Proposition 3.6] and [9, Proposition 6]):
Lemma 4.4.
Let $(M,g,f)$ be an $n$-dimensional complete weighted manifold. For $\kappa>0$, we assume that $\operatorname{Ric}_{f}^{1}\geq(n-1)\,\kappa\,\mathrm{e}^{-\frac{4f}{n-1}}\,g$ and $f$ is bounded from above. Let $\Sigma_{1}$ and $\Sigma_{2}$ be two complete properly immersed hypersurfaces in $M$ such that at least one of them is compact. Then we have
$d(\Sigma_{1},\Sigma_{2})\leq\frac{\mathrm{e}^{\frac{4\sup f}{n-1}}}{(n-1)\kappa}\left(\sup_{\Sigma_{1}}\left|\mathbf{H}_{f,\Sigma_{1}}\right|+\sup_{\Sigma_{2}}\left|\mathbf{H}_{f,\Sigma_{2}}\right|\right).$
Proof.
We may assume $\Sigma_{1}\cap\Sigma_{2}=\emptyset$. Set $d:=d(\Sigma_{1},\Sigma_{2})$. By the properness and compactness, there exists a unit speed minimal geodesic $\gamma:[0,d]\rightarrow M$ from $\Sigma_{1}$ to $\Sigma_{2}$. Notice that $\gamma$ intersects with $\Sigma_{1}$ and $\Sigma_{2}$ orthogonally by the first variation formula. By the same calculation as in the proof of [20, Proposition 3.6], it holds that
$0$ $\leq-\int_{0}^{d}\mathrm{e}^{\frac{2f(\gamma(t))}{n-1}}\operatorname{Ric}_{f}^{1}(\gamma^{\prime}(t),\gamma^{\prime}(t))\ \mathrm{d}t+\left(-\mathrm{e}^{\frac{2f(\gamma(d))}{n-1}}\,H_{f,\Sigma_{2}}(\gamma(d))+\mathrm{e}^{\frac{2f(\gamma(0))}{n-1}}\,H_{f,\Sigma_{1}}(\gamma(0))\right)$
$\leq-(n-1)\kappa\,\mathrm{e}^{-\frac{2\sup f}{n-1}}\,d+\mathrm{e}^{\frac{2\sup f}{n-1}}\left(\sup_{\Sigma_{2}}\left|\mathbf{H}_{f,\Sigma_{2}}\right|+\sup_{\Sigma_{1}}\left|\mathbf{H}_{f,\Sigma_{1}}\right|\right),$
where $H_{f,\Sigma_{1}}(\gamma(0))$ is the $f$-mean curvature for $\Sigma_{1}$ at $\gamma(0)$ with respect to $\gamma^{\prime}(0)$, and $H_{f,\Sigma_{2}}(\gamma(d))$ is that for $\Sigma_{2}$ at $\gamma(d)$ with respect to $\gamma^{\prime}(d)$. This proves the lemma. $\Box$
Lemma 4.4 leads us to the following intersection principle (cf. [9, Corollary 2]):
@@
Let $(M,g,f)$ be an $n$-dimensional complete weighted manifold. For $\kappa>0$, we assume that $\operatorname{Ric}_{f}^{1}\geq(n-1)\,\kappa\,\mathrm{e}^{-\frac{4f}{n-1}}\,g$ and $f$ is bounded from above. Then there exists a closed ball in $M$ such that every complete properly immersed $f$-minimal hypersurface in $M$ must intersect it.
Proof.
Fix $o\in M$ and $r>0$. Let $\Sigma$ be a complete properly immersed $f$-minimal hypersurface in $M$. From Lemma 4.4, it follows that
$d(\Sigma,\partial B_{r}(o))\leq\frac{\mathrm{e}^{\frac{4\sup f}{n-1}}}{(n-1)\kappa}\max_{\partial B_{r}(o)}\left|\mathbf{H}_{f,\partial B_{r}(o)}\right|.$
Notice that the right-hand side does not depend on $\Sigma$. Therefore, a closed ball centered at $o$ with radius large enough intersects with $\Sigma$. This completes the proof. $\Box$
The following is our main result (cf. [4, Theorem 1.3]):
Theorem 4.6.
For $3\leq n\leq 7$, let $(M,g,f)$ be an $n$-dimensional complete weighted manifold. For $\kappa>0$, we assume that $\operatorname{Ric}_{f}^{1}\geq(n-1)\,\kappa\,\mathrm{e}^{-\frac{4f}{n-1}}\,g$ and $f$ is bounded from above. For $\Lambda>0$ and a non-negative integer $I$, let $\mathcal{S}_{\Lambda,I}$ be the space of smooth complete embedded $f$-minimal hypersurfaces $\Sigma$ in $M$ with
@@
By Theorem 2.7, $M$ has finite fundamental group. By passing to the universal cover, we may assume that $M$ is simply connected. Notice that each hypersurface in $\mathcal{S}_{\Lambda,I}$ is properly embedded and two-sided (see [9, Propositions 3 and 8]). By virtue of [4, Proposition 3.3], every sequence $\{\Sigma_{i}\}_{i}$ in $\mathcal{S}_{\Lambda,I}$ has a convergent subsequence to a smooth complete properly embedded $f$-minimal hypersurface $\Sigma$ such that the convergence is smooth (possibly with finite multiplicity) away from a locally finite collection of points in the limit $\Sigma$. Note that $\Sigma$ is also non-trivial by Proposition 4.5. To show that the convergence is smooth everywhere, it suffices to verify that the multiplicity of the convergence is equal to one in view of the Allard regularity theorem. If it is greater than one, then $\Sigma$ is stable according to the argument of the proof of [4, Theorem 1.3] (cf. [12, Proposition 3.2], [9, Proposition 11]). This contradicts with Proposition 4.1, and hence the convergence is smooth everywhere. Also, by the same argument in [4, Theorem 1.3], we have $\operatorname{Ind}_{f}(\Sigma)\leq I$. Thus, we complete the proof. $\Box$
In a similar way, one can also obtain the following smooth compactness by using Theorem 2.7, Propositions 4.1 and 4.5 along the line of the proof of [9, Theorem 2]:
Theorem 4.7.
Let $(M,g,f)$ be a three dimensional complete weighted manifold. For $\kappa>0$, we assume that $\operatorname{Ric}_{f}^{1}\geq 2\,\kappa\,\mathrm{e}^{-2f}\,g$ and $f$ is bounded from above. For $\Lambda>0$ and a non-negative integer $\gamma$, let $\mathcal{S}_{\Lambda,\gamma}$ be the space of smooth complete embedded $f$-minimal surfaces $\Sigma$ in $M$ with
$m_{f,\Sigma}(\Sigma)\leq\Lambda,\quad\mathrm{genus}(\Sigma)\leq\gamma.$
Then $\mathcal{S}_{\Lambda,\gamma}$ is compact in the smooth topology.
Remark 4.8.
By slightly modifying the calculation in the proof of Lemma 4.4, the boundedness assumption for $f$ in Proposition 4.5, Theorems 4.6 and 4.7 can be weakened as the $f$-completeness in [56], [57].
Remark 4.9.
We give a remark on index and volume estimates in three dimensional case. Let $(M,g,f)$ be a three dimensional complete weighted manifold. Let $\Sigma$ stand for a closed embedded $f$-minimal surface in $M$ of genus $\gamma$. Let $U$ denote a bounded domain in $M$ with smooth boundary that contains $\Sigma$. We possess the following estimate of Ejiri-Micallef type (see [18], [4, Proposition 1.8]):
$\operatorname{Ind}_{f}(\Sigma)\leq C_{M,f,U}\left(m_{f,\Sigma}(\Sigma)+\gamma-1\right),$
where $C_{M,f,U}$ is a positive constant depending only on $M,f$ and $U$. Moreover, If we assume that $\operatorname{Ric}_{f}^{0}\geq K\,g$ for $K>0$, and the boundary of $U$ is convex, then
(4.3) $m_{f,\Sigma}(\Sigma)\leq\frac{16\pi}{K}(\gamma+1)\,\mathrm{e}^{-\inf_{\Sigma}f},$
which can be proved by using a first eigenvalue estimate of Choi-Wang type obtained in [20] along the line of the proof of [8, Proposition 8] (see [20, Theorem 3.1 and Remark 3.10]).
By Theorem 4.6 and Remark 4.9, we immediately obtain (cf. [37, Theorem 1]):
Corollary 4.10.
Let $(M,g,f)$ be a three dimensional closed weighted manifold of positive 0-weighted Ricci curvature. Then the space of smooth closed embedded $f$-minimal surfaces in $M$ of fixed topology is compact in the smooth topology.
We also have the following (cf. [8, Theorem 4]):
Corollary 4.11.
Let $(M,g,f)$ be a three dimensional complete weighted manifold. For $K>0$, we assume $\operatorname{Ric}_{f}^{0}\geq K\,g$. Let $U$ be a bounded domain in $M$. Let $\mathcal{S}_{\gamma,U}$ be the space of closed embedded $f$-minimal surfaces in $M$ with genus at most $\gamma$ and contained in the closure of $U$. If the closure of $U$ is contained in a bounded domain in $M$ with smooth convex boundary, then $\mathcal{S}_{\gamma,U}$ is compact in the smooth topology.
@@
Note that
$\operatorname{Ric}_{f}^{1}\geq\operatorname{Ric}_{f}^{0}\geq K\,g\geq(K\mathrm{e}^{2\,\inf f})\,\mathrm{e}^{-2f}\,g$
by (1.2), where the infimum is taken over the convex bounded domain. Hence, we arrive at the desired assertion by using Theorem 4.7 and (4.3) along the line of the second proof of [8, Theorem 4]. $\Box$
From Corollary 4.11, we conclude the following (cf. [8, Theorem 3]):
Corollary 4.12.
Let $(M,g,f)$ be a three dimensional complete weighted manifold. For $K>0$, we assume $\operatorname{Ric}_{f}^{0}\geq K\,g$. We also assume that $M$ admits an exhaustion by bounded domains with smooth convex boundary. For $D>0$ and a non-negative integer $\gamma$, let $\mathcal{S}_{D,\gamma}$ be the space of smooth closed embedded $f$-minimal surfaces $\Sigma$ in $M$ with
$\operatorname{diam}(\Sigma)\leq D,\quad\mathrm{genus}(\Sigma)\leq\gamma.$
Then $\mathcal{S}_{D,\gamma}$ is compact in the smooth topology.
@@
By the argument of the proof of [8, Lemma in p.364] together with Lemma 4.4, all surfaces in $\mathcal{S}_{D,\gamma}$ are contained in a closed ball in $M$. Thus, we complete the proof by applying Corollary 4.11 as in the proof of [8, Theorem 3]. $\Box$

### v2

@@ 4.1. Non-existence
Let $(M,g,f)$ be a complete weighted manifold of positive $1$-weighted Ricci curvature. Then there is no complete two-sided immersed stable $f$-minimal hypersurface $\Sigma$ in $M$ with $m_{f,\Sigma}(\Sigma)<+\infty$.
@@
Impera-Rimoldi [31] have presented non-existence results for stable $f$-minimal hypersurfaces satisfying volume growth conditions under the positivity of $\infty$-weighted Ricci curvature (see [31, Theorems A and B]). We show similar results for $1$-weighted Ricci curvature.
@@
The proof is done by contradiction. We possess the following relation of Brooks type between the bottom of the spectrum $\lambda_{1,\Sigma}$ of $\Delta_{f,\Sigma}$ and the volume entropy (cf. [7, Theorem 1], [31, Theorem A]):
@@ 2. $(2)$
We give a proof by contradiction. If $(1)$ holds, then [31, Proposition 6] together with Lemma 2.1 implies
@@
By the first condition in (4.2) and [31, Proposition 7], for every compact subset $\Omega$ in $\Sigma$, the bottom of the spectrum of the stability operator $\mathcal{L}_{f,\Sigma}$ over $\Sigma\setminus\Omega$ is negative. This contradicts with the finiteness of $f$-index and [31, Proposition 5]. $\Box$
@@ 4.2. Smooth compactness
Choi-Schoen [11] have established a smooth compactness of closed minimal surfaces in a compact three-dimensional manifold of positive Ricci curvature. Sharp [57] has extended their smooth compactness to the higher dimensional case. In the weighted case, such smooth compactness have been investigated by Li-Wei [40], Cheng-Mejia-Zhou [8], [9] and Barbosa-Sharp-Wei [4] under the positivity of $\infty$-weighted Ricci curvature (see also [13], [16] for self-shrinker). We prove smooth compactness under the positivity of $1$-weighted Ricci curvature. Our main result is the following (cf. [11], [40, Theorem 1]):
Theorem 4.4.
Let $(M,g,f)$ be a three dimensional closed weighted manifold of positive $1$-weighted Ricci curvature. Then the space of smooth closed embedded $f$-minimal surfaces in $M$ of fixed topology is compact in the smooth topology.
To prove it, we first present the following result of Sharp type (cf. [57], [4, Theorem 1.3]):
@@
For $3\leq n\leq 7$, let $(M,g,f)$ be an $n$-dimensional closed weighted manifold of positive $1$-weighted Ricci curvature. For $\Lambda>0$ and a non-negative integer $I$, let $\mathcal{S}_{\Lambda,I}$ be the space of smooth closed embedded $f$-minimal hypersurfaces $\Sigma$ in $M$ with
@@
By Proposition 2.7, the universal cover of $M$ is compact; in particular, $M$ has finite fundamental group (see also Remark 2.8). By passing to the universal cover, we may assume that $M$ is simply connected. Notice that each hypersurface in $\mathcal{S}_{\Lambda,I}$ is two-sided (cf. [9, Proposition 8]). With the help of [4, Proposition 3.3], every sequence $\{\Sigma_{i}\}_{i}$ in $\mathcal{S}_{\Lambda,I}$ has a convergent subsequence to a smooth closed embedded $f$-minimal hypersurface $\Sigma$ such that the convergence is smooth (possibly with finite multiplicity) away from a locally finite collection of points in the limit $\Sigma$. To show that the convergence is smooth everywhere, it suffices to verify that the multiplicity of the convergence is equal to one in view of the Allard regularity theorem. If it is greater than one, then $\Sigma$ is stable according to the argument of the proof of [4, Theorem 1.3] (cf. [13, Proposition 3.2], [9, Proposition 11]). This contradicts with Proposition 4.1, and hence the convergence is smooth everywhere. Also, by the same argument in [4, Theorem 1.3], we have $\operatorname{Ind}_{f}(\Sigma)\leq I$. Thus, we complete the proof. $\Box$
Due to Proposition 4.5, it suffices to derive index and volume estimates under the positivity of 1-weighted Ricci curvature in the three dimensional case. For the index, we possess the following estimate of Ejiri-Micallef type (see [19], [4, Proposition 1.8]):
Proposition 4.6 ([19], [4]).
Let $(M,g,f)$ be a three dimensional closed weighted manifold. Let $\Sigma$ be a closed embedded $f$-minimal surface in $M$ of genus $\gamma$. Then we have
$\operatorname{Ind}_{f}(\Sigma)\leq C_{M,f}\left(m_{f,\Sigma}(\Sigma)+\gamma-1\right).$
We now aim to obtain volume estimates. We deduce such estimates from a lower bound of the first eigenvalue of the Laplacian induced from an affine connection. On an $n$-dimensional weighted manifold $(M,g,f)$, we consider an affine connection
(4.3) $\mathsf{D}_{X}Y:=\nabla_{X}Y-\frac{\mathrm{d}f(X)}{n-1}Y-\frac{\mathrm{d}f(Y)}{n-1}X.$
It has been observed that the Ricci tensor induced from this affine connection coincides with $\operatorname{Ric}^{1}_{f}$ (see [61, Proposition 3.3], [41]). Li-Xia [41] have studied a family of affine connections including (4.3), and obtained an integral formula of Reilly type. In a special case for (4.3), it can be stated as follows (see [41, Theorem 1.1], and also [22, Proposition 2.5, Remark 2.6]):
Proposition 4.7 ([41]).
Let $(M,g,f)$ be an $n$-dimensional compact weighted manifold with boundary. Then for every $\varphi\in C^{\infty}(M)$,
$\quad\,\,\int_{M}\left\{\left(\Delta_{\frac{n}{n-1}f}\,\varphi\right)^{2}-\operatorname{Ric}_{f}^{1}\left(\nabla\varphi,\nabla\varphi\right)-\left|\mathrm{Hess}\,\varphi-\frac{1}{n-1}\langle\nabla\varphi,\nabla f\rangle g\right|^{2}\right\}\ \mathrm{d}m_{f}$
$=\int_{\partial M}\left(\varphi_{\nu}^{\,2}H_{f,\partial M}+2\varphi_{\nu}\Delta_{f,\partial M}\,\psi+\mathrm{II}_{\partial M}\left(\nabla_{\partial M}\,\psi,\nabla_{\partial M}\,\psi\right)\right)\ \mathrm{d}m_{f,\partial M},$
where $\psi:=\varphi|_{\partial M}$, and $\nu$ is the outer unit normal vector field on $\partial M$.
In order to formulate the integral formula for general affine connections, Li-Xia [41] have introduced the notion of the affine Laplacian (see [41, Definition 2.4]). On a closed embedded hypersurface $\Sigma$ in an $n$-dimensional closed weighted manifold $(M,g,f)$, the affine Laplacian associated with (4.3) is given by
$\Delta^{\mathsf{D}}_{\Sigma}:=\mathrm{e}^{\frac{f}{n-1}}\,\Delta_{f,\Sigma}.$
By direct calculations, it coincides with the weighted Laplacian over a weighted manifold
$\left(\Sigma,\mathrm{e}^{-\frac{f}{n-1}}\,g_{\Sigma},\frac{n+1}{2(n-1)}f\right).$
In particular, it is self-adjoint with respect to the weighted measure $m_{\frac{n}{n-1}f,\Sigma}$, and its first non-zero eigenvalue has the following variational characterization:
(4.4) $\lambda^{\mathsf{D}}_{1,\Sigma}=\inf_{\phi}\,\frac{\displaystyle\int_{\Sigma}\,|\nabla_{\Sigma}\,\phi|^{2}\,\mathrm{d}m_{f,\Sigma}}{\displaystyle\int_{\Sigma}\,\phi^{2}\,\mathrm{d}m_{\frac{n}{n-1}f,\Sigma}},$
where the infimum is taken over all $\phi\in C^{\infty}(\Sigma)\setminus\{0\}$ with $\int_{\Sigma}\,\phi\,\mathrm{d}m_{\frac{n}{n-1}f,\Sigma}=0$. Furthermore, we consider a conformally deformed metric
(4.5) $g_{f}:=\mathrm{e}^{-\frac{2n}{(n-1)^{2}}f}g$
over $M$. Notice that its volume measure $v_{g_{f},\Sigma}$ over $\Sigma$ coincides with $m_{\frac{n}{n-1}f,\Sigma}$, and therefore
(4.6) $\lambda^{\mathsf{D}}_{1,\Sigma}=\inf_{\phi}\,\frac{\displaystyle\int_{\Sigma}\,|\nabla_{\Sigma}\,\phi|^{2}_{g_{f}}\,\mathrm{e}^{-\frac{n+1}{(n-1)^{2}}f}\,\mathrm{d}v_{g_{f},\Sigma}}{\displaystyle\int_{\Sigma}\,\phi^{2}\,\mathrm{d}v_{g_{f},\Sigma}}\leq\mathrm{e}^{-\frac{n+1}{(n-1)^{2}}\inf_{\Sigma}f}\lambda^{g_{f}}_{1,\Sigma}$
by (4.4), where $\lambda^{g_{f}}_{1,\Sigma}$ is the first non-zero eigenvalue of the Laplacian on $\Sigma$ induced from $g_{f}$.
We obtain the following eigenvalue estimate of Choi-Wang type (cf. [12], [8, Theorem 2], [40, Theorem 7], [45, Theorem 3], [21, Theorem 1.1]):
Proposition 4.8.
Let $(M,g,f)$ be an $n$-dimensional closed weighted manifold. For $K>0$, we assume $\operatorname{Ric}_{f}^{1}\geq K\,\mathrm{e}^{-\frac{f}{n-1}}\,g$. Let $\Sigma$ be a closed embedded $f$-minimal hypersurface in $M$. Then we have
(4.7) $\lambda^{\mathsf{D}}_{1,\Sigma}\geq\frac{K}{2}.$
@@
The desired estimate has been proved by the first named author [21] for general affine connections (see [21, Theorem 1.1]). For the completeness, we here give its proof. Proposition 2.7 tells us that the universal cover of $M$ is compact; in particular, $M$ has finite fundamental group (see also Remark 2.8).
We first show the estimate when $M$ is simply connected. Then $M$ and $\Sigma$ are orientable since $\Sigma$ is closed and embedded. Further, by virtue of Proposition 2.10, $\Sigma$ must be connected. Therefore, it divides $M$ into two components $U_{1}$ and $U_{2}$ (see e.g., [40, Lemma 6]). Let $u$ be an eigenfunction for $\lambda^{\mathsf{D}}_{1,\Sigma}$. We may assume
$\int_{\partial U_{1}}{\rm I\hskip-0.09995ptI}_{\partial U_{1}}(\nabla_{\partial U_{1}}\,u,\nabla_{\partial U_{1}}\,u)\ \mathrm{d}m_{f,\partial U_{1}}\geq 0.$
Let $\varphi$ be the solution to the following boundary value problem:
$\begin{cases}\Delta_{\frac{n}{n-1}f}\,\varphi=0&\mbox{ on }U_{1},\\ \varphi=u&\mbox{ on }\partial U_{1}.\end{cases}$
By using Proposition 4.7 over $U_{1}$, we obtain
$0$ $\geq K\int_{U_{1}}|\nabla\varphi|^{2}\ \mathrm{d}m_{\frac{n}{n-1}f}-2\lambda^{\mathsf{D}}_{1,\Sigma}\int_{\partial U_{1}}\varphi_{\nu}\,u\ \mathrm{d}m_{\frac{n}{n-1}f,\partial U_{1}}+\int_{\partial U_{1}}{\rm I\hskip-0.09995ptI}_{\partial U_{1}}(\nabla_{\partial U_{1}}\,u,\nabla_{\partial U_{1}}\,u)\ \mathrm{d}m_{f,\partial U_{1}}$
$\geq(K-2\lambda^{\mathsf{D}}_{1,\Sigma})\int_{U_{1}}|\nabla\varphi|^{2}\ \mathrm{d}m_{\frac{n}{n-1}f},$
where $\nu$ denotes the outer unit normal vector field on $\partial U_{1}$. This leads us to (4.7).
For the general case, in view of the finiteness of the fundamental group, the lift $\overline{\Sigma}$ of $\Sigma$ is a closed embedded $\bar{f}$-minimal hypersurface in the universal cover, where $\bar{f}$ is the lift of $f$. The result in the above paragraph says that the first non-zero eigenvalue $\lambda^{\mathsf{D}}_{1,\overline{\Sigma}}$ of $\Delta^{\mathsf{D}}_{\overline{\Sigma}}$ satisfies
$\lambda^{\mathsf{D}}_{1,\Sigma}\geq\lambda^{\mathsf{D}}_{1,\overline{\Sigma}}\geq\frac{K}{2}.$
This completes the proof. $\Box$
We arrive at the following volume estimate of Yang-Yau type (cf. [62], [8, Proposition 8]):
Proposition 4.9.
Let $(M,g,f)$ be a three dimensional closed weighted manifold. For $K>0$, we assume $\operatorname{Ric}_{f}^{1}\geq K\,\mathrm{e}^{-\frac{f}{2}}\,g$. Let $\Sigma$ be a closed embedded $f$-minimal hypersurface in $M$. Then
$m_{f,\Sigma}(\Sigma)\leq\frac{16\pi}{K}\left(\frac{2}{|\pi_{1}(M)|}-\frac{1}{2}\chi(\Sigma)\right)\,\mathrm{e}^{-\inf_{\Sigma}f+\frac{1}{2}\sup_{\Sigma}f},$
where $\chi(\Sigma)$ denotes the Euler characteristic of $\Sigma$.
@@
Similarly to the proof of Proposition 4.8, we first prove the estimate when $M$ is simply connected. Then $\Sigma$ is orientable, and we denote its genus by $\gamma$. We consider the conformally deformed metric $g_{f}$ defined as (4.5). Similarly to the proof of [8, Proposition 7], applying the classical Yang-Yau estimate to the metric $g_{f}$ tells us that
(4.8) $\mathrm{e}^{-\frac{1}{2}\sup_{\Sigma}f}m_{f,\Sigma}(\Sigma)\leq m_{\frac{3}{2}f,\Sigma}(\Sigma)=v_{g_{f},\Sigma}(\Sigma)\leq\frac{8\pi(\gamma+1)}{\lambda^{g_{f}}_{1,\Sigma}}.$
On the other hand, by combining (4.6) and Proposition 4.8, we possess
(4.9) $\mathrm{e}^{-\inf_{\Sigma}f}\lambda^{g_{f}}_{1,\Sigma}\geq\lambda^{\mathsf{D}}_{1,\Sigma}\geq\frac{K}{2}.$
From (4.8) and (4.9), it follows that
(4.10) $\mathrm{e}^{-\frac{1}{2}\sup_{\Sigma}f}m_{f,\Sigma}(\Sigma)\leq\frac{16\pi}{K}(\gamma+1)\,\mathrm{e}^{-\inf_{\Sigma}f}.$
This proves the desired one. For the general case, we complete the proof by applying (4.10) to the universal cover similarly to the proof of [8, Proposition 8]. $\Box$
We are now in a position to conclude Theorem 4.4.
Proof of Theorem 4.4.
It immediately follows from Propositions 4.5, 4.6 and 4.9. $\Box$